OpenStudy (zyberg):

Father says to his son: "When I will be as old as my dad is now, you will be 5 times older than right now, and your age will be 8 years older than I am right now. My father and mine ages sum is 100 years." Find son's age.

OpenStudy (zyberg):

My attempt to solve this: D - grandfather's age; d - father's age; s - son's age. Then, I tried to make a system of equations: D = 5s, (D - d) + s = d + 8, D + d = 100. However, the solution for son that I find from this is 12. Then grandfathers age would be 12 * 5 = 60, however with that fathers age would need to be 40 and that doesn't make any sense with the second equation... Please help me to spot my mistakes!

OpenStudy (jiteshmeghwal9):

Let father's age be x & that of son's be y & that of grandfather's age be z. Also let the difference between grandfather's age & father 's age be 't'. x+y=100 z-x=t Now when father will achieve the grandfather's age the son will be y+t years old or we can say that his age will be y+8 i.e; 8 years older than father's age y+t=x+8 y+t=5y

OpenStudy (zyberg):

@jiteshmeghwal9 Solved your system of equations. x + y = 100 should be x + z = 100 tho. Another thing to notice that we get grandfather's age to be less than fathers... x = 78; z = 22

OpenStudy (jiteshmeghwal9):

yeah tht was a typo x+z=100 z-x=t when father will achieve the grandfather's age , t years would have been passed from present time x+8=y+t y+t=5y

OpenStudy (zyberg):

@jiteshmeghwal9 as I said, I get that grandfather is thrice as old as the father, using your equations.

OpenStudy (jiteshmeghwal9):

but yeah i was putting my opinion here I m not trying to confuse u

OpenStudy (zyberg):

Oh, thank you! Seems that I had made a mistake while solving them. @ganeshie8, I think that the solution is right. Logically, it seems to make sense?

ganeshie8 (ganeshie8):

Yeah, jitesh's solution looks good to me. I was hoping to get an integer answer... question appears to be wrong.

OpenStudy (zyberg):

Well, perhaps they tried to trick people into thinking that it is wrong to get non integer solutions... although I was hoping for that as well, but...